# On the best approximation by finite Gaussian mixtures

### v1

@@ Appendix D Proof of the convergence rates of NPMLE
The following lemma bounds the $L_{\infty}$-metric entropy of the Gaussian mixture densities with compactly supported and sub-Weibull mixing distributions. Note that this is an improvement of the previous result [Zha09, Lemma 2] which deals with truncated $L_{\infty}$-norm on a compact interval. Combined with our main result in Theorem 1, this also improves the untruncated $L_{\infty}$-entropy estimate in [GvdV07b, Lemma 2] which gives $\log N(\epsilon,{\mathcal{F}}_{{\mathcal{P}}^{\textsf{Bdd}}_{M}},L_{\infty})\lesssim M\log\frac{1}{\epsilon}\log\frac{M}{\epsilon}$. The key idea is to use the smoothness of the Gaussian mixtures to relate $L_{\infty}$-error (pointwise) to $L_{1}$-error (TV), the latter of which can be further bounded by moment matching.
@@
$|{\mathcal{C}}|\leq\binom{m+|{\mathcal{L}}|-1}{m}|{\mathcal{N}}_{m}|\overset{\rm(a)}{\leq}\frac{\left(\frac{2t}{\epsilon}+m\right)^{m}}{m!}2m\left(1+\frac{1}{\epsilon}\right)^{m-1}\leq 2m\left(\frac{2t}{\epsilon}+m\right)^{m}\left(\frac{2e}{\epsilon m}\right)^{m},$
@@
$\log N(2\epsilon,{\mathcal{F}}_{\mathcal{P}},\mathsf{TV})\leq m\log\left(\frac{4et}{\epsilon^{2}m}+\frac{2e}{\epsilon}\right)+\log 2m\overset{\rm(a)}{\leq}O\left(m\log\frac{1}{\epsilon}\right),$ (61)
@@
The proof for ${\mathcal{P}}={\mathcal{P}}_{\alpha}(\beta)$ is analogous, where we need to choose $t$ depending on $m$ and $\beta$ as in the proof of Theorem 4. Specifically, let $t=t(m,\beta)\asymp_{\alpha}m^{\frac{2}{2+\alpha}}\beta^{\frac{\alpha}{2+\alpha}}$ if $m^{\alpha-2}\lesssim_{\alpha}\beta^{2\alpha},$ and $t\asymp_{\alpha}\beta\left(m\log\frac{m^{\alpha-2}}{\beta^{2\alpha}}\right)^{\frac{1}{\alpha}}$ otherwise. Plugging $t=t(m,\beta)$ into (61) and noting that $m\gtrsim_{\alpha}\beta$ by Theorem 7, we likewise obtain that $\log{\mathcal{N}}(\epsilon,{\mathcal{F}}_{\mathcal{P}},L_{\infty})\lesssim_{\alpha}\left(m\log\frac{1}{\epsilon}\right)$. ∎

### v2

@@ Appendix D Proof of the convergence rates of NPMLE
D.1 Unconstrained NPMLE
Lemma 12.
Let ${\mathcal{Q}}_{0}$ be the collection of all distributions on $\mathbb{R}$. Then, for any $\epsilon\in(0,1/2]$,
$m^{\star}(\epsilon,{\mathcal{Q}}_{0},L_{\infty,[-M,M]})\lesssim M\sqrt{\log\epsilon^{-1}}\vee\log\epsilon^{-1}.$
Proof.
Set $t=M+4\sqrt{\log{\frac{1}{\epsilon}}}$ and $m=12t\sqrt{\kappa\log\frac{1}{\epsilon}}$, where $\kappa$ is defined in Theorem 3. Fix any $P\in{\mathcal{Q}}_{0}$, and define $P_{I_{t}}$ and $P_{I_{t}^{c}}$ as $P$ conditioned on $I_{t}=[-t,t]$ and $I_{t}^{c}$, respectively. Then, we have
$f_{P}=P(I_{t})f_{P_{I_{t}}}+P(I_{t}^{c})f_{P_{I_{t}^{c}}}.$
Applying Theorem 3 with $3\sqrt{\kappa}t\leq m\leq\kappa t^{2}$ yields that ${\mathcal{E}}^{\star}(m,{\mathcal{P}}^{\textsf{Bdd}}_{t},\chi^{2})\leq\epsilon^{8}$. By Lemma 4, there exists $\tilde{P}\in{\mathcal{P}}_{m}$ such that $\mathsf{TV}(f_{\tilde{P}},f_{P_{I_{t}}})\leq\epsilon^{4}$. Since $\|f_{P}^{\prime}\|_{\infty}\leq\|\phi^{\prime}\|_{\infty}=(2\pi e)^{-1/2}$, $\mathsf{TV}(f_{\tilde{P}},f_{P_{I_{t}}})\leq\epsilon^{4}$ then implies $\|{f_{\tilde{P}}-f_{P_{I_{t}}}}\|_{\infty}\leq(\frac{2}{\pi e})^{1/4}\epsilon^{2}$. Let $P_{m+1}=P(I_{t})\tilde{P}+P(I_{t}^{c})\delta_{t}\in{\mathcal{P}}_{m+1}$. Then, we have
$|f_{P}(x)-f_{P_{m+1}}(x)|$ $=\left|P(I_{t})f_{P_{I_{t}}}(x)+P(I_{t}^{c})f_{P_{I_{t}^{c}}}(x)-P(I_{t})f_{\tilde{P}}(x)-P(I_{t}^{c})\phi(x-t)\right|$
$\leq\left|f_{P_{I_{t}}}(x)-f_{\tilde{P}}(x)\right|+\left|f_{P_{I_{t}^{c}}}(x)\right|+\left|\phi(x-t)\right|.$
Since $P_{I_{t}^{c}}$ is supported on $\{|x|\geq t\}$, we have $f_{P_{I_{t}^{c}}}(x)\leq\phi(t-M)\leq\epsilon^{8}$ for any $|x|\leq M$. Hence,
$\sup_{|x|\leq M}|f_{P}(x)-f_{P_{m+1}}(x)|$ $\leq\left(\frac{2}{\pi e}\right)^{1/4}\epsilon^{2}+2\epsilon^{8}\leq\epsilon.$
Consequently, $m^{\star}(\epsilon,{\mathcal{Q}}_{0},L_{\infty,[-M,M]})\leq m+1$, and the desired result follows. ∎
D.2 Constrained NPMLE
@@
The following lemma bounds the $L_{\infty}$-metric entropy of the Gaussian mixture densities with compactly supported and sub-Weibull mixing distributions. Note that this is an improvement of the previous result [Zha09, Lemma 2] which deals with truncated $L_{\infty}$-norm on a compact interval. Combined with our main result in Theorem 1, this also improves the untruncated $L_{\infty}$-entropy estimate in [GvdV07, Lemma 2] which gives $\log N(\epsilon,{\mathcal{F}}_{{\mathcal{P}}^{\textsf{Bdd}}_{M}},L_{\infty})\lesssim M\log\frac{1}{\epsilon}\log\frac{M}{\epsilon}$. The key idea is to use the smoothness of the Gaussian mixtures to relate $L_{\infty}$-error (pointwise) to $L_{1}$-error (TV), the latter of which can be further bounded by moment matching.
@@
$|{\mathcal{C}}|\leq\binom{m+|{\mathcal{L}}|-1}{m}|{\mathcal{N}}_{m}|\overset{\mathrm{(a)}}{\leq}\frac{\left(\frac{2t}{\epsilon}+m\right)^{m}}{m!}2m\left(1+\frac{1}{\epsilon}\right)^{m-1}\leq 2m\left(\frac{2t}{\epsilon}+m\right)^{m}\left(\frac{2e}{\epsilon m}\right)^{m},$
@@
$\log N(2\epsilon,{\mathcal{F}}_{\mathcal{P}},\mathsf{TV})\leq m\log\left(\frac{4et}{\epsilon^{2}m}+\frac{2e}{\epsilon}\right)+\log 2m\overset{\mathrm{(a)}}{\leq}O\left(m\log\frac{1}{\epsilon}\right),$ (68)
@@
The proof for ${\mathcal{P}}={\mathcal{P}}_{\alpha}(\beta)$ is analogous, where we need to choose $t$ depending on $m$ and $\beta$ as in the proof of Theorem 4. Specifically, let $t=t(m,\beta)\asymp_{\alpha}m^{\frac{2}{2+\alpha}}\beta^{\frac{\alpha}{2+\alpha}}$ if $m^{\alpha-2}\lesssim_{\alpha}\beta^{2\alpha},$ and $t\asymp_{\alpha}\beta\left(m\log\frac{m^{\alpha-2}}{\beta^{2\alpha}}\right)^{\frac{1}{\alpha}}$ otherwise. Plugging $t=t(m,\beta)$ into (68) and noting that $m\gtrsim_{\alpha}\beta$ by Theorem 6, we likewise obtain that $\log{\mathcal{N}}(\epsilon,{\mathcal{F}}_{\mathcal{P}},L_{\infty})\lesssim_{\alpha}{m\log\frac{1}{\epsilon}}$. ∎